OpenStudy (anonymous):

Given sinA=3/5 and cosB=-1/4, A in quadrant II and B in quadrant III. Find exact value of tan (A-B)

OpenStudy (bittuaryan):

@Meinme Kindly show the steps

OpenStudy (anonymous):

@bittuaryan sorry i misread the question

OpenStudy (bittuaryan):

Bravo....u can do....try again

OpenStudy (anonymous):

 this is what i got using formula of tan a+b and ASTC rule

OpenStudy (bittuaryan):

if u got: \[\frac{ (\frac{ -5 }{ 4 } )+\sqrt{15}}{ 1+5(\frac{ \sqrt{15} }{ 8 })}\] =0.76681639399